# Entanglement and Bell inequality violation in vector diboson systems produced in decays of spin-0 particles

### v1

@@ Abstract
We discuss entanglement and the violation of the CGLMP inequality in a system of two vector bosons produced in the decay of a spin-0 particle. We assume the most general CPT conserving, Lorentz- invariant coupling of the spin-0 particle with the daughter bosons. We compute the most general two-boson density matrix obtained by averaging over kinematical configurations with an appropriate probability distribution (which can be obtained when both bosons subsequently decay into fermion-antifermion). We show that the two-boson state is entangled and violates the CGLMP inequality for all values of the (anomalous) coupling constants and that in this case the state is entangled iff it can violate the CGLMP inequality. As an exemplary process of this kind we use the decay $H\to ZZ$ with anomalous coupling.
@@ I Introduction
One of the most intriguing and fascinating aspects of quantum mechanics is quantum nonlocality. Nonlocality is expressed by correlations of entangled quantum states of quantum objects. As was proven theoretically by John Stuart Bell [1] nonlocality is an immanent aspect of quantum reality which was confirmed for elementary systems in series of correlation experiments in different scales of energy-momentum as well as different conditions for space-time localization [2, 3, 4, 5, 6, 7, 8, 9]. Recently, the possibility of testing nonlocality of quantum mechanics under extremal conditions arising in a decay of the Higgs boson has been put forward [10, 11, 12, 13, 14, 15, 16, 17]. Indeed, in the most interesting case of Higgs decay into two gauge bosons ($ZZ$ or $WW$) we have possibility to measure quantum spin states of unstable vector bosons. This is possible by means of registration of their decay products (leptons) which are detected in polarization states determined by kinematics. The decay process of weak bosons takes place in the extremely short time $10^{-25}$ sec and at last one of decaying bosons is in a virtual state, out of the mass shell. Therefore, investigation of possible correlations of entangled vector bosons can provide a test of quantum mechanical nonlocality under completely new conditions.
@@
Einstein–Podolsky–Rosen (EPR) type correlation experiment with relativistic vector bosons in the simplest scalar state was considered in [18] for the first time. The possibility of observing the violation of Clauser–Horn–Shimony–Holt (CHSH) and Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequalities by $WW$ bosons arising in the actual decay $H\to WW$ for the first time was analyzed in [10]. In [11] the possibility of violation of CHSH, Mermin and CGLMP inequalities by a boson-antiboson system in the most general scalar state (introduced in this context in [19]) was discussed. In [12] the entanglement and violation of CGLMP inequality by a pair of $Z$ bosons arising in the decay $H\to ZZ$ was considered. The authors of [12] assumed the Standard Model interaction of $H$ with the daughter $Z$ bosons and showed that $ZZ$ state produced in such a process is highly entanglement. In our paper [16] we also discussed entanglement and the violation of CGLMP inequality by a $ZZ$ pair arising in the decay $H\to ZZ$ but assuming anomalous (beyond the Standard Model) structure of the vertex describing interaction of a Higgs particle with two daughter bosons. It can be shown (compare e.g. [20, 21]) that the amplitude corresponding to the most general Lorentz-invariant, CPT conserving coupling of the pseudoscalar/scalar particle $X$ with two vector bosons $V_{1}$, $V_{2}$ in the decay
@@
It is worth noticing that the Higgs decay is not the only process which was proposed as a test bed for exploring fundamental quantum properties like entanglement or Bell-type inequality violation in high energy physics. In fact, the first system considered in this context was a system of top quarks produced in colliders [23, 24, 25, 26, 27, 28, 29, 30]. Moreover, the only experimental observation of entanglement at such high energy scale has been recently reported by the ATLAS collaboration at the LHC in a $tt$ system [31] and subsequently confirmed by the CMS collaboration [32].
Other high energy processes have also been proposed in this context, including, among other: various scattering processes [33, 34], $B^{0}\bar{B}^{0}$ mesons [35], $tW$ systems [36], $WW$ pairs produced in electron-positron colliders [37], $\tau\tau$ pairs [38] and tripartite systems [39, 40].
@@ II Decay of a pseudoscalar/scalar particle into two vector bosons
Next, without loss of generality we can assume that bosons arising in the decay (1) move along $z$-axis, i.e. we can take $^{\mu}=(\omega_{1},\textbf{k})$ k ), $p^{\mu}=(\omega_{1},-\textbf{k})$, where $\textbf{k}=(0,0,|\textbf{k}|)$ and energies $\omega_{1}$, $\omega_{2}$ are given explicitly in (58,59). We also simplify the notation of basis two-boson states in this case
@@
$\rho(m_{1},m_{2},c,\tilde{c})\\ =|\psi_{VV}^{\mathsf{norm}}(m_{1},m_{2},c,\tilde{c})\rangle\langle\psi_{VV}^{\mathsf{norm}}(m_{1},m_{2},c,\tilde{c})|.$ (20)
@@
where $\mathcal{P}_{c,\tilde{c}}(m_{1},m_{2})$ is a normalized probability distribution. The explicit form of this probability distribution can be determined in the case when the daughter bosons arising in (1) subsequently decay into massless fermions [20, 16]
$X\to VV\to f_{1}^{+}f_{1}^{-}f_{2}^{+}f_{2}^{-}.$ (22)
Following the same line of reasoning as in our previous paper [16] we find
@@
$\mathsf{f}$ $=-(c+1)\mathsf{B}(1)+c\mathsf{B}(-1)$
$\phantom{=-(c+1)}+2i\tilde{c}\big{[}(1+c)\tilde{\mathsf{B}}(0)-c\tilde{\mathsf{B}}(-2)\big{]},$ (30)
@@ IV Violation of Bell inequalities
$(O_{A}\otimes I)\mathcal{O}_{\mathsf{Bell}}(U_{V}(0,0),U_{V}(\tfrac{\pi}{2},0),U_{V}(t,\theta),U_{V}(-t,\theta))\\ (O_{A}\otimes I)$ (46)
@@
$\sin(\alpha+\theta)=0,\qquad\cos(\alpha+\theta)=\pm 1,$ (49)
$\cos t=\tfrac{a}{\sqrt{a^{2}+4|a_{13}|^{2}}},$ (50)
$\sin t=\mp\tfrac{2|a_{13}|}{\sqrt{a^{2}+4|a_{13}|^{2}}},$ (51)
@@
In our exemplary decay $H\to ZZ$ from Eqs. (3235) we see that $a_{13}\not=0$. Therefore, in this decay the CGLMP inequality can be violated for all values of $c$, $\tilde{c}$. In Fig. 2 we have plotted the value of $({\mathcal{I}}_{3})_{\mathrm{max}}$ obtained with the help of the above optimization procedure for that decay.
@@
Experimentally, the state $\rho_{ZZ}(c,\tilde{c})$ is reconstructed in collider experiments via quantum tomography methods [13, 52]. In such a case the presence of errors and background in the process $H\to ZZ\to f_{1}^{+}f_{1}^{-}f_{2}^{+}f_{2}^{-}$ modifies the state (27). To estimate how this modification influences the violation of the CGLMP inequality we consider the resistance of this violation with respect to the white noise. The noise resistance we define as a minimal value of $\lambda$, $\lambda_{\mathsf{min}}$, for which the state
@@ V Conclusions
It is very interesting that, as we have shown, the state $\rho_{VV}(c,\tilde{c})$ (27) violates the CGLMP inequality iff it is entangled. It is a non-trivial observation since for an arbitrary $3\times 3$ quantum state $\rho$ such a statement is true only if $\rho$ is pure.
This also leads to strong phenomenological implications, since we have proven that in a pair of vector bosons one can indirectly test the Bell-type inequality violation by checking that the pair is entangled. And this is a much easier experimental task. Moreover, this observation holds regardless of the interaction among the spin-0 particle and the gauge bosons (provided that they are CPT conserving and Lorentz invariant, a very sensible requirement), so no future new physics can change the statement of entanglement and Bell violation in this system.

### v2

@@ Abstract
We discuss entanglement and the violation of the Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequality in a system of two vector bosons produced in the decay of a spin-0 particle. We assume the most general CPT conserving, Lorentz- invariant coupling of the spin-0 particle with the daughter bosons. We compute the most general two-boson density matrix obtained by averaging over kinematical configurations with an appropriate probability distribution (which can be obtained when both bosons subsequently decay into fermion-antifermion). We show that the two-boson state is entangled and violates the CGLMP inequality for all values of the (anomalous) coupling constants and that in this case the state is entangled iff it can violate the CGLMP inequality. As an exemplary process of this kind we use the decay $H\to ZZ$ with anomalous coupling.
@@ I Introduction
One of the most intriguing and fascinating aspects of quantum mechanics is quantum nonlocality. Nonlocality is expressed by correlations of entangled quantum states of quantum objects. As was proven theoretically by John Stuart Bell [1] nonlocality is an immanent aspect of quantum reality which was confirmed for elementary systems in series of correlation experiments in different scales of energy-momentum as well as different conditions for space-time localization [2, 3, 4, 5, 6, 7, 8, 9]. Recently, the possibility of testing Bell-type inequality violation under extremal conditions arising in a decay of the Higgs boson has been put forward [10, 11, 12, 13, 14, 15, 16, 17]. Indeed, in the most interesting case of Higgs decay into two gauge bosons ($ZZ$ or $WW$) we have possibility to measure quantum spin states of unstable vector bosons. This is possible by means of registration of their decay products (leptons) which are detected in polarization states determined by kinematics. The decay process of weak bosons takes place in the extremely short time $10^{-25}$ sec and at last one of decaying bosons is in a virtual state, out of the mass shell. Therefore, investigation of possible correlations of entangled vector bosons can provide a test of quantum mechanics under completely new conditions.
It should be stressed here that analysis of spin correlations in the relativistic regime is hindered by problems with the definition of a proper spin observable for a relativistic particle. Various operators have been proposed in the literature, see, e.g., [18, 19, 20, 21, 22, 23] and references therein. In our paper we used the so called Newton-Wigner spin operator corresponding to a spin of a particle in its rest frame, we justified this choice in, e.g., [11]. We will also shortly comment this point in Conclusions.
@@
Einstein–Podolsky–Rosen (EPR) type correlation experiment with relativistic vector bosons in the simplest scalar state was considered in [24] for the first time. The possibility of observing the violation of Clauser–Horn–Shimony–Holt (CHSH) and Collins–Gisin–Linden–Massar–Popescu (CGLMP) inequalities by $WW$ bosons arising in the actual decay $H\to WW$ for the first time was analyzed in [10] (notice, that the first proper discussion of spin correlations in this decay was given in [25]). In [11] the possibility of violation of CHSH, Mermin and CGLMP inequalities by a boson-antiboson system in the most general scalar state (introduced in this context in [26]) was discussed. In [12] the entanglement and violation of CGLMP inequality by a pair of $Z$ bosons arising in the decay $H\to ZZ$ was considered. The authors of [12] assumed the Standard Model interaction of $H$ with the daughter $Z$ bosons and showed that $ZZ$ state produced in such a process is highly entanglement. In our paper [16] we also discussed entanglement and the violation of CGLMP inequality by a $ZZ$ pair arising in the decay $H\to ZZ$ but assuming anomalous (beyond the Standard Model) structure of the vertex describing interaction of a Higgs particle with two daughter bosons. It can be shown (compare e.g. [27, 28]) that the amplitude corresponding to the most general Lorentz-invariant, CPT conserving coupling of the pseudoscalar/scalar particle $X$ with two vector bosons $V_{1}$, $V_{2}$ in the decay
@@
It is worth noticing that the Higgs decay is not the only process which was proposed as a test bed for exploring fundamental quantum properties like entanglement or Bell-type inequality violation in high energy physics. In fact, the first propositions of EPR-like experiments in particle physics were put forward more then forty years ago [30, 31]. In recent years, the first system considered in this context was a system of top quarks produced in colliders [32, 33, 34, 35, 36, 37, 38, 39]. Moreover, the only experimental observation of entanglement at such high energy scale has been recently reported by the ATLAS collaboration at the LHC in a $tt$ system [40] and subsequently confirmed by the CMS collaboration [41].
Other high energy processes have also been proposed in this context, including, among other: various scattering processes [42, 43], $B^{0}\bar{B}^{0}$ mesons [44], $tW$ systems [45], $WW$ pairs produced in electron-positron colliders [46], $\tau\tau$ pairs [47] and tripartite systems [48, 49, 50].
@@ II Decay of a pseudoscalar/scalar particle into two vector bosons
We do not provide a Lagrangian of the interaction, since the anomalous vertices are obtained from operators carrying a dimension larger than four. Thus, it would be necessary to collect all dimension-six operators in the SMEFT (Standard Model Effective Field Theory) associated with this interaction which is out of the scope of the present work.
@@
Next, without loss of generality we can assume that bosons arising in the decay (1) move along $z$-axis, i.e. we can take $^{\mu}=(\omega_{1},\textbf{k})$ k ), $p^{\mu}=(\omega_{1},-\textbf{k})$, where $\textbf{k}=(0,0,|\textbf{k}|)$ and energies $\omega_{1}$, $\omega_{2}$ are given explicitly in (56,57). We also simplify the notation of basis two-boson states in this case
@@
$\rho(m_{1},m_{2},c,\tilde{c})=|\psi_{VV}^{\mathsf{norm}}(m_{1},m_{2},c,\tilde{c})\rangle\langle\psi_{VV}^{\mathsf{norm}}(m_{1},m_{2},c,\tilde{c})|.$ (20)
@@
where $\mathcal{P}_{c,\tilde{c}}(m_{1},m_{2})$ is a normalized probability distribution. The explicit form of this probability distribution can be determined for different channels of the subsequent decay of daughter vector bosons $VV$ arising in (1). In our exemplary case of the Higgs decay into Z bosons, the decay chain $H\to ZZ^{*}\to(f^{+}f^{-})(f^{+}f^{-})$ constitutes one of the most promising channels to certify entanglement at colliders in a qutrit-qutrit system. Thus we stick to this channel in our general analysis. In the case when the daughter bosons decay into massless fermions
$X\to VV\to f_{1}^{+}f_{1}^{-}f_{2}^{+}f_{2}^{-},$ (22)
we can use the results from [27, 16] and following exactly the same line of reasoning as in our previous paper (compare Eqs. (33–38) from [16]) we find
@@
$\mathsf{f}$ $=-(c+1)\mathsf{B}(1)+c\mathsf{B}(-1)+2i\tilde{c}\big{[}(1+c)\tilde{\mathsf{B}}(0)-c\tilde{\mathsf{B}}(-2)\big{]},$ (30)
@@ IV Violation of Bell inequalities
$(O_{A}\otimes I)\mathcal{O}_{\mathsf{Bell}}(U_{V}(0,0),U_{V}(\tfrac{\pi}{2},0),U_{V}(t,\theta),U_{V}(-t,\theta))(O_{A}\otimes I)$ (46)
@@
$\sin(\alpha+\theta)=0,\qquad\cos(\alpha+\theta)=\pm 1,\\ \cos t=\tfrac{a}{\sqrt{a^{2}+4|a_{13}|^{2}}},\qquad\sin t=\mp\tfrac{2|a_{13}|}{\sqrt{a^{2}+4|a_{13}|^{2}}},$ (49)
@@
In our exemplary decay $H\to ZZ$ from Eqs. (3235) we see that $a_{13}\not=0$. Therefore, in this decay the CGLMP inequality can be violated for all values of $c$, $\tilde{c}$. In Fig. 2 we have plotted the value of $({\mathcal{I}}_{3})_{\mathrm{max}}$ obtained with the help of the above optimization procedure for that decay. In Fig. 3 we have plotted $({\mathcal{I}}_{3})_{\mathrm{max}}$ as a function of $c$ for three chosen values of $\tilde{c}=0,0.5,1.5$. The value $\tilde{c}=0$ corresponds to the scalar Higgs, $\tilde{c}=0.5$ corresponds to the boundary value given in Eq. (64) while $\tilde{c}=1.5$ is given for comparison.
@@
Figure 3: In this figure we present the maximal value of $\mathcal{I}_{3}$ in the state (27) as a function of $c$, for three chosen values of $\tilde{c}$. We have inserted the measured values for the Higgs mass, Z mass and Z decay width, i.e., we put $\mathsf{b}_{Z}$, $\mathsf{e}_{Z}$, $\mathsf{f}_{Z}$ and $\mathsf{h}_{Z}$ given in Eqs. (3235). Notice that the choice $\tilde{c}=0$ corresponds to the Standard Model value, $\tilde{c}=0.5$ to the experimental bound (64). The value $\tilde{c}=1.5$ is presented for comparison.
@@
Experimentally, the state $\rho_{ZZ}(c,\tilde{c})$ is reconstructed in collider experiments via quantum tomography methods [13, 62]. In such a case the presence of errors and background in the process $H\to ZZ\to f_{1}^{+}f_{1}^{-}f_{2}^{+}f_{2}^{-}$ modifies the state (27). To estimate how this modification influences the violation of the CGLMP inequality we consider the resistance of this violation with respect to the white noise. It is worth noticing that the addition of white noise can also model, in the first approximation, decoherence of the final state. The noise resistance we define as a minimal value of $\lambda$, $\lambda_{\mathsf{min}}$, for which the state
@@ V Conclusions
In Introduction we noticed that different spin operators have been proposed in the literature. In this context it is interesting to mention the paper [22] in which spin correlations of an electron-positron pair arising in a decay of a pseudo-scalar particle were discussed. The authors of [22] use different spin operator from the operator used in the present work. As a result they obtain that the considered $e^{+}e^{-}$ pair becames unentangled in the high energy limit.
It is also worth noticing here that spin correlation functions in colliders are determined from cross-sections (or angular distribution of momenta). This allows for construction of peculiar local hidden variable models duplicating experimental statistics ([64, 65]). One can hope that this loophole will be closed with future technical developments as it was the case in the standard low-energy Bell tests [66, 7, 67]. The problem of loophole-free tests of Bell nonlocality at colliders has been recently addressed in [68]. The authors of this paper take a different point of view. They point out that by performing the quantum tomography of the system we are able to reconstruct the full two-partite density matrix. Then, any realization of the observables $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ over the density matrix can be theoretically performed.
In this context it is very interesting that, as we have shown, the state $\rho_{VV}(c,\tilde{c})$ (27) violates the CGLMP inequality iff it is entangled. It is a non-trivial observation since for an arbitrary $3\times 3$ quantum state $\rho$ such a statement is true only if $\rho$ is pure. This also leads to strong phenomenological implications, since we have proven that in a pair of vector bosons one can indirectly test the Bell-type inequality violation by checking that the pair is entangled. And this seems to be a much easier experimental task (compare the recent paper [40]). Moreover, this observation holds regardless of the interaction among the spin-0 particle and the gauge bosons (provided that they are CPT conserving and Lorentz invariant, a very sensible requirement), so no future new physics can change the statement of entanglement and Bell violation in this system.
It is also important to stress that this relation between entanglement and the violation of the CGLMP inequality is solely based on the texture of the matrix (which is a consequence of the symmetries involved in the decay) and not on the experimental way of getting the density matrix itself.
@@
Finally, In our paper as an exemplary process we considered the decay $H\to ZZ$. However, some scalar or pseudo-scalar mesons could also decay in a similar way. For scalar mesons $\tilde{c}=0$ but one could have $c\not=0$ while for pseudo-scalar ones $\tilde{c}\not=0$.